# Federated Learning: A new frontier in the exploration of multi-institutional medical imaging data

### v1

@@ 6.4 Security
Selective parameter sharing: Privacy can also be violated by overfitting a locally trained model. The selective sharing of parameters is a solution to handle this problem. As an example, the Privacy-preserving Federated Brain Tumor Segmentation [Li et al., 2019] selectively shares local information, that is, model components $\Delta{\boldsymbol{\Theta}}_{c}$ whose absolute values are greater than the chosen threshold are transferred to the server. In addition, elements of $\Delta{\boldsymbol{\Theta}}_{c}$ are clipped to a specific range of $[-\gamma,\gamma]$. This method has been shown to be effective in segmenting brain tumors.
@@ 7 Federated learning frameworks
APPFL (Advanced Privacy-Preserving Federated Learning) [Li et al., 2024] is a high-performance framework and benchmarking tool that handles various communication protocols, data transfer and compression methods, and privacy preservation strategies. APPFL supports the Message Passing Interface standard for single-machine and cluster servers and gRPC/Globus Compute for distributed training.
@@
FedTree [Li et al., 2023b] is a specialized tool for FL systems using tree-based models. It is designed to be highly efficient, effective, and secure. It delivers a range of features, including parallel computing on multi-core CPUs and GPUs, support for HE, secure aggregation, and DP.
@@ 8 Discussion
We have been observing an enormous interest in FL over the past decade, resulting in a full gamut of new aggregation and learning techniques, as previously mentioned in sections 5 and 6. However, those methods often only partially address the problems or, not surprisingly, create new ones that require an in-depth examination. In this section, we thoroughly discuss these issues before MI can fully leverage the benefits of FL.
@@ 8.1 Model accuracy
We start our discussion with some important comments about the accuracy of the global model. Once training the global model is initiated, one has to ensure that the involved sites do not negatively affect the aggregation function [Darzidehkalani et al., 2022]. Usually, this is achieved by dynamically assigning a weighting factor to each client based on its local loss function [Shen et al., 2021, Mächler et al., 2022, 2023, Wicaksana et al., 2023, Wu et al., 2023b, Liu et al., 2023a] or accuracy-related metrics [Yang et al., 2023]. The weighting factor may also cover the similarity between the parameters of transferred local and global models [Khan et al., 2022, 2023, Yang et al., 2023]. Another approach is to abandon specific parameters of the local models during the aggregation process. The parameters to be withdrawn might be randomly selected [Gunesli et al., 2021] or chosen from specific layers (e.g., normalization layer) [Bernecker et al., 2022].
To prevent global model deterioration, it is fundamental to avoid forgetting previously learned knowledge. One can achieve this through the curriculum learning [Jiménez-Sánchez et al., 2023], where inconsistent predictions caused by forgotten data are penalized. An alternative approach is continuous learning, which introduces and manages new memories without forgetting the previously gained knowledge [Sheller et al., 2020, Huang et al., 2022]. The consequence of catastrophic interference is even more critical as the training might introduce a bias towards the most recent clients involved in the training procedure. To handle the bias favoring new clients, one can use the cyclic weight transfer approach, which trains the model at the client and then transfers model parameters to the next one [Darzidehkalani et al., 2022]. Dynamic weight correction is another technique that stabilizes the training process and avoids a drift between the global model and local models [Yang et al., 2022b].
@@ 8.2 Data heterogeneity
The most commonly used method to aggregate MI data is the FedAvg [McMahan et al., 2017]. However, the FedAvg-derived results strongly depend on the homogeneity of the data between the sites. If the data is heterogeneous, the accuracy of the global model decreases, and the model performs satisfactorily only for some clients [Zhao et al., 2018]. However, although the FedProx is a typical option for MI [Li et al., 2020a], it does not work adequately in all scenarios and therefore does not necessarily always show significantly better results than the FedAvg [Luo and Wu, 2022, Zhou et al., 2023, Yang et al., 2023]. To ensure the best possible results, participants, before training starts, must find a consensus on the criteria for selecting the best possible global model, such as the metric chosen at a satisfactory level or the number of global rounds that must be passed [Darzidehkalani et al., 2022].
@@
Another possible solution to handle data heterogeneity is to harmonize the data or standardize the pre-processing pipeline across the collaborators [Jiang et al., 2022]. These may require sharing metadata between the institutions or even exemplary acquisitions to set up a harmonization protocol. Precisely, data harmonization protocols typically require selecting a target institution to which the data is normalized, and hence, sharing exemplary acquisitions with other clients is favorable. This might be challenging to achieve, considering the privacy constraints. Therefore, the cooperation of clinicians and technicians is essential to standardize the way data is handled in different institutions [Darzidehkalani et al., 2022]. It is also beneficial to create model datasets that can serve as a target site for harmonization and to establish universal metrics used to determine the quality of harmonized data [Pinto et al., 2020].
Yet another practical approach to deal with data heterogeneity is direct [Guo et al., 2021] or indirect [Liu et al., 2021a, Luo and Wu, 2022, Zhao and Huang, 2023] data sharing (see Fig. 11). This strategy addresses the problem from a statistical point of view by sharing the distribution of classes in local data or a small amount of data that contains examples from each class. It enables clients to augment their datasets and create an independently and identically distributed (IID) dataset.
@@ 8.4 Security
The most often exploited methods to preserve data privacy are DP, secure MPC, and HE. The first method, the DP, introduces random noise to the dataset to reduce the risk of revealing sensitive personal information. The commonly used algorithm for training local models is the DP-SGD [Ziller et al., 2021b, Malekzadeh et al., 2021, Nguyen et al., 2022a, Kalra et al., 2023]. However, this technique may also reduce the global model accuracy [Truong et al., 2021, Abad et al., 2022]. In contrast to DP, the MPC preserves high global model accuracy. It enables multiple sites to compute a function with their inputs without revealing them to other parties [Mothukuri et al., 2021, Truong et al., 2021]. HE is yet another technique that can improve data privacy and is often used for secure aggregation [Malekzadeh et al., 2021, Makkar and Santosh, 2023]. Here, the computations are performed on encrypted data, and the results can only be decrypted by the institution that requests it [Fereidooni et al., 2021, Truong et al., 2021]. However, the MPC and HE increase the computational burden and complicate the training algorithm, consequently significantly increasing the local training duration [Mothukuri et al., 2021, Truong et al., 2021].
@@
Due to the high prerequisites of privacy protection techniques, it is essential to consider whether it is necessary or not to implement them in the FL-based system. Profitability depends largely on the level of trust between the parties involved and the scale of the project [Darzidehkalani et al., 2022]. If the institutions trust each other or the training involves a small group of clients who can sign appropriate data protection agreements, employing the above methods might be unfounded. Nevertheless, this does not exempt the clients from protecting their data from other parties, i.e., clarifying the required level of trust among clients is paramount before introducing a federated learning architecture [Kandati and Gadekallu, 2023a].

### v2

@@ 6.4 Security
Selective parameter sharing: Privacy can also be violated by overfitting a locally trained model. The selective sharing of parameters is a solution to handle this problem. As an example, the Privacy-preserving Federated Brain Tumor Segmentation [Li et al., 2019] selectively shares local information, that is, model components $\Delta{\boldsymbol{\Theta}}_{c}$ whose absolute values are greater than the chosen threshold are transferred to the server. In addition, elements of $\Delta{\boldsymbol{\Theta}}_{c}$ are clipped to a specific range of $\gamma,\gamma]$.
6.5 Model heterogeneity
Model heterogeneity can significantly restrict the application of classic FL in medical centers. This limitation arises because it requires all clients to use the same neural network architecture, which is often not feasible due to varying computational resources or differences in the data they handle. One potential solution to this challenge is model-agnostic FL, which imposes no specific requirements on client models. This approach typically relies on knowledge distillation or generative methods.
Knowledge distillation: Federated Similarity Knowledge Distillation (FedSKD) [Weng et al., 2025] is an aggregation-free method and works in a peer-to-peer manner, enabling bidirectional knowledge exchange among clients through multi-dimensional similarity knowledge distillation (SKD). FedSKD categorizes models into two types: 1) local models, $\mathbf{\Theta}_{c}$, which adapt to local data distribution, and 2) models received from other participants, $\tilde{\mathbf{\Theta}}_{c^{\prime}}$, that carry cross-client knowledge. In each global round, a client either trains its model locally or receives a model from another institution, selected in a predetermined order, and then performs mutual distillation. The mutual distillation process has two stages. First, domain-specific knowledge injection, in which the local model distils its expertise into the received model using SKD. Second, cross-client knowledge absorption, in which the received model enhances the local model’s generalisation capabilities. These steps are repeated until the parameters of all clients’ models converge. The total loss of SKD training for the $r$-th round on a client can be defined as follows:
$\mathcal{L}(\mathbf{\Theta}_{c}^{r},\tilde{\mathbf{\Theta}}_{c^{\prime}}^{r};\mathcal{D}_{c})={}\mathcal{L}_{\mathrm{CE}}(\mathbf{\Theta}_{c}^{r};\mathcal{D}_{c})+\mathcal{L}_{\mathrm{CE}}(\tilde{\mathbf{\Theta}}_{c^{\prime}}^{r};\mathcal{D}_{c})\\ +\lambda\mathcal{L}_{\mathrm{SKD}}(\mathbf{\Theta}_{c}^{r},\tilde{\mathbf{\Theta}}_{c^{\prime}}^{r};\mathcal{D}_{c}),$ (38)
where $\mathcal{L}_{\mathrm{CE}}$ and $\mathcal{L}_{\mathrm{SKD}}$ are the cross-entropy loss and multi-dimensional SKD loss, respectively, and $\lambda$ is a balancing hyperparameter.
Generative methods: An example of a generative method is the Federated Generative Autoregressive Transformer (FedGAT) [Nezhad et al., 2025]. This approach facilitates collaboration among clients using different reconstruction model architectures through a two-tier strategy. In the first tier, a GAT prior captures the distribution of MRI data from multiple sites. The GAT prior includes a variational autoencoder that encodes MRI images into discrete token maps at various spatial scales, which are then modeled by an autoregressive transformer to predict feature maps at higher scales. In the second tier, each institution trains its own preferred reconstruction model on local data and fine-tunes on a hybrid dataset generated by augmenting local MRI data with synthetic images from the global GAT prior.
@@ 7 Federated learning frameworks
APPFL (Advanced Privacy-Preserving Federated Learning) [Li et al., 2025c] is a high-performance framework and benchmarking tool that handles various communication protocols, data transfer and compression methods, and privacy preservation strategies. APPFL supports the Message Passing Interface standard for single-machine and cluster servers and gRPC/Globus Compute for distributed training.
@@
Fed-BioMed [Cremonesi et al., 2025] is an open-source research and development initiative aimed at translating collaborative learning, including FL, into practical medical applications. It offers a variety of features, including a demonstrated framework for deploying FL within hospital networks and user-friendly tools for data management and client participation. Additionally, Fed-BioMed facilitates the easy implementation of state-of-the-art methods while ensuring compliance with data providers’ privacy and governance requirements.
FedTree [Li et al., 2023a] is a specialized tool for FL systems using tree-based models. It is designed to be highly efficient, effective, and secure. It delivers a range of features, including parallel computing on multi-core CPUs and GPUs, support for HE, secure aggregation, and DP.
@@
8 Real-world implementations of federated learning in medical imaging
The application of FL in MI is predominantly limited to simulations rather than real-world deployments. In these scenarios, data from a single source is typically experimentally divided to mimic the local datasets of individual clients [Malekzadeh et al., 2021, Stripelis et al., 2021b, Islam et al., 2022, Gupta et al., 2025]. Alternatively, data from various institutions or publicly available datasets is handled to simulate multi-sites [Jiang et al., 2023, Dalmaz et al., 2024, Tan et al., 2024, Wang et al., 2024]. However, the clients in these simulations are not authentic medical infrastructures – they are virtual nodes or independent computing processes operating on a centralised research cluster. This means that the federation is simulated in a controlled manner, bypassing the complexities of real clinical systems.
Recently, several real-world implementations of FL in clinical settings have appeared. These developments provide an opportunity for a more thorough analysis and understanding of the practical challenges associated with implementing FL in clinical environments, particularly those related to the General Data Protection Regulation (GDPR) and the technical aspects of hospital infrastructure.
8.1 Clinical cases of federated learning
Prediction of COVID-19 clinical outcomes: Dayan et al. [2021] Dayan et al. [2021] developed EXAM, a global FL model to predict the future oxygen requirements of symptomatic COVID-19 patients. The model combines two types of data: 1) chest X-ray images and 2) Electronic Medical Record (EMR) data. The FL architecture employs a client-server architecture, with the model implemented in TensorFlow (https://www.tensorflow.org/) using the NVIDIA Clara Train SDK (https://docs.nvidia.com/clara/index.html). This project involved 20 institutions across four continents. Due to variations in device manufacturers, imaging protocols and demographics, the datasets were heterogeneous, concerning both X-ray images and EMR feature distributions. The final EXAM model achieved an average Area Under the Curve (AUC) greater than 0.92, indicating a 16% improvement in AUC and a 38% increase in generalizability compared to models trained locally at individual sites, with significant benefits for institutions with small or unbalanced datasets.
Rare cancer boundary detection: Pati et al. [2022] implemented an FL framework to detect the sub-compartment boundaries of glioblastoma using multi-parametric MRI scans. The FL architecture features a central aggregation server at the University of Pennsylvania, securely connected to all collaborating sites. Communication during the training process is protected using the Transport Layer Security cryptographic protocol, while datasets are aggregated within Trusted Execution Environments using the Intel Software Guard Extensions (SGX; https://www.intel.com/content/www/us/en/developer/tools/software-guard-extensions/overview.html) for confidentiality and integrity. This project includes 71 sites across six continents, with a standardised data preprocessing pipeline that accounts for variations in scanner hardware and acquisition protocols. The final consensus model demonstrated significant improvements over the initially publicly trained model, with 33% improvement in delineating the surgically targetable tumour and 23% improvement in defining the complete tumour extent.
Lung pathology segmentation: Bujotzek et al. [2025] demonstrated a proof-of-concept training of segmentation models for lung pathology detection using CT data in an FL environment. The datasets contained a variety of lung pathologies and exhibited significant differences in annotation procedures, scanner manufacturers, and acquisition protocols. The results consistently demonstrated that FL approaches outperformed local models in all evaluation scenarios. The experiment was conducted across six university hospitals in Germany: Charité Berlin, Technical University of Munich, University Medicine Essen, and the hospitals in Frankfurt am Main, Cologne, and Kiel.
Breast cancer classification: Tzortzis et al. [2025] implemented an end-to-end FL framework for breast cancer classification, specifically targeting BIRADS classification (Breast Imaging Reporting and Data System) with mammography data. This project infrastructure, developed under the INCISIVE European Project (https://incisive-project.eu/), consists of a central cluster and local clusters at hospitals. Clients operate on-site using the infrastructure that meets specified minimum requirements. Built on Kubernetes, the framework uses a standard API for communication, enabling centralized control and simplifying maintenance for hospital staff. This study involved three hospital institutions: the Hellenic Cancer Society in Greece, the University of Novi Sad in Serbia, and Aristotle University of Thessaloniki in Greece. The datasets were heterogeneous and contained artefacts, including breast implants, surgical clips, and annotation labels. As a result, the model performance improved by approximately 35% at the first participating hospital, aligning the FL-based results with those of a centralized model.
Cardiac CT imaging: Tölle et al. [2025] implemented an FL system to enhance the analysis of cardiac CT imaging for patients undergoing transcatheter aortic valve implantation. The project uses a centralized FL approach with a central server located at Heidelberg University. Different institutions used the same physical machine for training, although their network configurations varies. This study involved eight cardiology and radiology departments, all affiliated with the German Center for Cardiovascular Diseases (DZHK). Datasets varied in sample count and acquisition protocols, with only a small fraction labelled, necessitating a semi-supervised federated knowledge distillation strategy. The resulting model showed improved generalizability in downstream tasks, such as coronary artery segmentation, indicating a better understanding of the global context.
8.2 Challenges in deploying federated learning in a clinical scenario
Implementing FL-based infrastructure in a real clinical practice demonstrated numerous organisational, technical, and legal challenges, requiring significant effort and specific mitigation strategies.
Data heterogeneity has emerged as the major challenge for all studies attempting to deploy FL in clinic settings. This issue arises from differences in patient demographics, variations in imaging protocols and acquisition equipment (such as different scanner manufacturers) [Tzortzis et al., 2025, Tölle et al., 2025]. Additionally, data quality inconsistencies and a lack of data acquisition standardization across institutions contribute to this problem [Dayan et al., 2021, Pati et al., 2022, Bujotzek et al., 2025]. As a result, these factors lead to decreased accuracy and limit the network’s generalizability. Tzortzis et al. [2025] observed that data heterogeneity significantly reduced the performance of the global model at one hospital, noting a nearly 40% decrease in predictive performance, as explained by the F1 score, compared to the standard deep learning. In most cases, the authors sought to minimise this problem by focusing on data preparation, such as detailed data specifications [Bujotzek et al., 2025], targeted pre-processing procedures [Tölle et al., 2025] or data harmonization [Tzortzis et al., 2025]. The inability to access the raw training data directly complicated the debugging procedure for unsatisfactory outcomes [Dayan et al., 2021, Pati et al., 2022]. As a result, the authors recommended appointing a Data Manager or a Curator at each institution to handle local dataset curation and ensure compliance with data-related requirements [Tzortzis et al., 2025].
Another significant challenge involves the infrastructure and logistics. Implementing an FL-based system often requires installing a specialised platform within highly restricted clinical information technology environments, enabling local model training while maintaining secure, compliant communication with a central server Bujotzek et al. [2025]. The authors therefore emphasise the importance of designing frameworks that are easy to deploy, potentially leveraging tools such as Kubernetes, and that are sufficiently intuitive for routine use by hospital staff [Tzortzis et al., 2025]. Other identified challenges include disparities in institutional resources, including hardware capabilities and network bandwidth, on-site computational power and storage management, and the integration with third-party systems, such as PACS and annotation tools [Bujotzek et al., 2025, Tölle et al., 2025]. To address some of these issues, Tölle et al. [2025] standardized the infrastructure by providing identical hardware to all participating institutions. However, differences in network configurations and software installations across sites continued to create difficulties. Infrastructure instability is also a challenge, with sites dropping out due to issues such as straggling, crashes, and disconnected nodes. These issues require experiment restarts, leading to idle machines at other locations [Bujotzek et al., 2025]. Asynchronous FL is therefore recommended to effectively manage nodes with heterogeneous resources [Tzortzis et al., 2025].
Security, confidentiality, and legal compliance have posed significant barriers to implementing FL in clinical settings. Healthcare applications, especially in the European Union and the United States, must adhere to strict regulations such as the GDPR and the Health Insurance Portability and Accountability Act (HIPAA) [Pati et al., 2022, Tzortzis et al., 2025]. While the FL principle generally aligns well with data protection laws, FL-based systems are not automatically compliant with the GDPR or similar regulations. Simply keeping data locally does not provide sufficient privacy warranties [Truong et al., 2021]. Besides, the model parameters shared among participants can retain sensitive information that may be exploited in advanced attacks, such as model inversion [Truong et al., 2021, Tölle et al., 2025, Tzortzis et al., 2025]. To address these issues in practice, collaborative networks use signed data-sharing agreements among trusted partners [Tzortzis et al., 2025]. Technical mitigations include employing Transport Layer Security, username/password authentication, and server-side IP address whitelisting to ensure secure communication [Tölle et al., 2025]. Security-enhancing features, such as partial weight sharing, to minimise the risk of information leakage from model gradients, have also been examined [Dayan et al., 2021]. Likewise, clients are required to perform privacy impact assessments before joining FL networks [Tzortzis et al., 2025].
Figure 14: Issue–method–effect diagram illustrating how different aggregation strategies (top graph) and learning methods (bottom graph) address key challenges in federated learning. The color-encoded lines indicate the relationships between specific issues, the methods used to mitigate them, and the resulting benefits. Methods that address multiple challenges are highlighted with multiple colors and connections.
@@ 9 Discussion
We have been observing an enormous interest in FL over the past decade, resulting in a full gamut of new aggregation and learning techniques, as previously mentioned in sections 5 and 6. We analyzed distributions of these methods, categorized them by imaging modalities and deep learning tasks (see Fig. 5). These graphs lead to some interesting conclusions. 1) The standard FedAvg and FL-EV are leading the way to aggregate the models, accounting for more than half of medical applications across modalities and tasks. 2) The same aggregation schemes have been used in many contexts, which highlights their universal applicability, although it should be noted that not every algorithm might be suitable for every task and/or modality. 3) Concerning the learning procedures employed in the federated learning system, aside from fully supervised training, the most frequently used strategies across modalities and tasks are split learning and knowledge distillation. 4) Interestingly, the usability of various techniques may be (potentially) limited in some particular applications. For instance, in a translation/synthesis task, fully supervised training and split learning have been used so far.
Figure 14 provides a concise overview of aggregation strategies and learning methods, highlighting the problems they aim to solve and summarizing their main effects. However, these approaches often only partially address the problems or, not surprisingly, create new ones that require an in-depth examination. In this section, we thoroughly discuss these issues before MI can fully leverage the benefits of FL.
@@ 9.1 Model accuracy
We start our discussion with some important comments about the accuracy of the global model. Once training the global model is initiated, one has to ensure that the involved sites do not negatively affect the aggregation function [Darzidehkalani et al., 2022b]. Usually, this is achieved by dynamically assigning a weighting factor to each client based on its local loss function [Shen et al., 2021, Mächler et al., 2022, 2023, Wicaksana et al., 2023, Wu et al., 2024, Liu et al., 2023a] or accuracy-related metrics [Yang et al., 2023]. The weighting factor may also cover the similarity between the parameters of transferred local and global models [Khan et al., 2022, 2023, Yang et al., 2023]. Another approach is to abandon specific parameters of the local models during the aggregation process. The parameters to be withdrawn might be randomly selected [Gunesli et al., 2021] or chosen from specific layers (e.g., normalization layer) [Bernecker et al., 2022].
To prevent global model deterioration, it is fundamental to avoid forgetting previously learned knowledge. One can achieve this through the curriculum learning [Jiménez-Sánchez et al., 2023], where inconsistent predictions caused by forgotten data are penalized. An alternative approach is continuous learning, which introduces and manages new memories without forgetting the previously gained knowledge [Sheller et al., 2020, Huang et al., 2022]. The consequence of catastrophic interference is even more critical as the training might introduce a bias towards the most recent clients involved in the training procedure. To handle the bias favoring new clients, one can use the cyclic weight transfer approach, which trains the model at the client and then transfers model parameters to the next one [Darzidehkalani et al., 2022b]. Dynamic weight correction is another technique that stabilizes the training process and avoids a drift between the global model and local models [Yang et al., 2022b].
@@ 9.2 Data heterogeneity
The most commonly used method to aggregate MI data is the FedAvg [McMahan et al., 2017]. However, the FedAvg-derived results strongly depend on the homogeneity of the data between the sites. If the data is heterogeneous, the accuracy of the global model decreases, and the model performs satisfactorily only for some clients [Zhao et al., 2018]. However, although the FedProx is a typical option for MI [Li et al., 2020a], it does not work adequately in all scenarios and therefore does not necessarily always show significantly better results than the FedAvg [Luo and Wu, 2022, Zhou et al., 2023, Yang et al., 2023]. FedBN [Li et al., 2021] is often considered a baseline strategy for handling data heterogeneity by excluding the parameters of batch normalisation layers during aggregation. This method can reduce distributional shifts across clients. However, its effectiveness may vary depending on the amount of local data available, as batch normalisation statistics are computed from local samples. To ensure the best possible results, participants, before training starts, must find a consensus on the criteria for selecting the best possible global model, such as the metric chosen at a satisfactory level or the number of global rounds that must be passed [Darzidehkalani et al., 2022b].
@@
Another possible solution to handle data heterogeneity is to harmonize the data or standardize the pre-processing pipeline across the collaborators [Jiang et al., 2022]. These may require sharing metadata between the institutions or even exemplary acquisitions to set up a harmonization protocol. Precisely, data harmonization protocols typically require selecting a target institution to which the data is normalized, and hence, sharing exemplary acquisitions with other clients is favorable. This might be challenging to achieve, considering the privacy constraints. Therefore, the cooperation of clinicians and technicians is essential to standardize the way data is handled in different institutions [Darzidehkalani et al., 2022b]. It is also beneficial to create model datasets that can serve as a target site for harmonization and to establish universal metrics used to determine the quality of harmonized data [Pinto et al., 2020].
Yet another practical approach to deal with data heterogeneity is direct [Guo et al., 2021] or indirect [Liu et al., 2021a, Luo and Wu, 2022, Zhao and Huang, 2023, Li et al., 2025a] data sharing (see Fig. 12). This strategy addresses the problem from a statistical point of view by sharing the distribution of classes in local data or a small amount of data that contains examples from each class. It enables clients to augment their datasets and create an independently and identically distributed (IID) dataset.
@@ 9.4 Security
The most often exploited methods to preserve data privacy are DP, secure MPC, and HE. The first method, the DP, introduces random noise to the dataset to reduce the risk of revealing sensitive personal information. The commonly used algorithm for training local models is the DP-SGD [Ziller et al., 2021b, Malekzadeh et al., 2021, Nguyen et al., 2022a, Kalra et al., 2023]. However, this technique may also reduce the global model accuracy [Truong et al., 2021, Abad et al., 2022]. The trade-off between security and neural network performance vastly depends on the model architecture and the privacy regime. According to studies by Ziller et al. [2021b] and Nguyen et al. [2022a], the reduction in accuracy typically ranges from about 5 to 15 percentage points. Such a decrease may be unacceptable for models intended for clinical practice. In contrast to DP, the MPC preserves high global model accuracy. It enables multiple sites to compute a function with their inputs without revealing them to other parties [Mothukuri et al., 2021, Truong et al., 2021]. HE is yet another technique that can improve data privacy and is often used for secure aggregation [Malekzadeh et al., 2021, Makkar and Santosh, 2023]. Here, the computations are performed on encrypted data, and the results can only be decrypted by the institution that requests it [Fereidooni et al., 2021, Truong et al., 2021]. However, the MPC and HE increase the computational burden and complicate the training algorithm, consequently significantly increasing the local training duration [Mothukuri et al., 2021, Truong et al., 2021]. Research on the application of HE in FL for MI shown that the time required for joint model training can increase up to six times compared to FL without HE [Lessage et al., 2024]. Additionally, even small changes to the model architecture, such as adding a single layer, can increase inference time by up to eight times when HE is used [Dutil et al., 2021].
@@
Due to the high prerequisites of privacy protection techniques, it is essential to consider whether it is necessary or not to implement them in the FL-based system. Profitability depends largely on the level of trust between the parties involved and the scale of the project [Darzidehkalani et al., 2022b]. If the institutions trust each other or the training involves a small group of clients who can sign appropriate data protection agreements, employing the above methods might be unfounded. Nevertheless, this does not exempt the clients from protecting their data from other parties, i.e., clarifying the required level of trust among clients is paramount before introducing a federated learning architecture [Kandati and Gadekallu, 2023a].